6. Check whether -150 is a term of the A.P. 11, 8, 5, 2, …

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
For the given series, A.P. 11, 8, 5, 2.. First term, a = 11 Common difference, d = a2−a1 = 8−11 = −3 Let −150 be the nth term of this A.P. As we know, for an A.P., an = a+(n−1)d -150 = 11+(n -1)(-3) -150 = 11-3n +3 -164 = -3n n = 164/3 Clearly, n is not an integer but a fraction. Therefore, – 150 is not a term of this A.P.